# RL of Thoughts: Navigating LLM Reasoning with Inference-time Reinforcement Learning

### v2

@@ Appendix I Analyses on the Design of Logic Blocks
To examine the impact of each block, we remove each of them and train a navigator:
Table 17: Ablation study results for each logic block.
LLM Ablation MATH GPQA StrategyQA Average
Qwen2.5-Instruct-7B w/o Decompose 75.42 31.92 77.00 61.45
w/o Debate 74.02 36.61 77.58 62.74
w/o Refine 75.76 41.29 72.93 63.33
Full RLoT 76.70 44.64 79.04 66.79
GPT-4o-mini w/o Decompose 76.08 41.74 79.33 65.72
w/o Debate 74.68 45.31 78.75 66.25
w/o Refine 76.44 51.56 74.38 67.46
Full RLoT 77.36 54.02 82.68 71.35
The ablation results in Table 17 confirm that all blocks are effective.
@@ 3 Logical blocks (action space)
For instance, "Decompose" reflects the widely studied divide-and-conquer approach, while "Debate" aligns with strategies involving comparative evaluation (Wang & Chiew, 2010). Also, they are empirically effective. Prior studies (Madaan et al., 2023; Xue et al., 2024) have demonstrated that "Decompose" and "Refine" can significantly enhance the performance of LLMs on complex tasks.
@@
Third, here is an input-output example for the ’Debate’ block (as Appendix B):
@@ 18 C2: Completeness within the step:
5Please conduct the following Subtask{subtask_id} to continue the reasoning. text_decompose.

### v3

@@ I.1 Effect of Logic Block Number
I.2 Role of Different Logic Blocks on Different Datasets
@@ 3 Logical blocks (action space)
For instance, "Decompose" reflects the widely studied divide-and-conquer approach, while "Debate" aligns with strategies involving comparative evaluation (Wang and Chiew, 2010). Also, they are empirically effective. Prior studies (Madaan et al., 2023; Xue et al., 2024) have demonstrated that "Decompose" and "Refine" can significantly enhance the performance of LLMs on complex tasks.
@@
Third, here is an input-output example for the ’Debate’ block (as Appendix B.1):
@@ 18 C2: Completeness within the step:
5Please conduct the following Subtask{subtask_id} to continue the reasoning.
@@
K.3 Sensitive to action-prompts phrasing.
To evaluate whether the performance of RLoT is sensitive to the specific wording of the action prompts, we conducted a robustness test by rephrasing the natural language instructions for the Decompose, Debate, and Refine actions.
Decompose: We rephrase the prompts for the action of Decompose as:
Listing 11: Prompts for obtaining subtasks in action “Decompose”
1This is the problem and the reasoning progress so far.
2{Problem and previous steps}
3Break down the current problem into smaller, manageable sub-problems.
4Solving these sub-problems sequentially should lead to the solution of the original question.
5Only provide subtasks decomposition here, and DONOT conduct specific reasoning or calculation.
6Format: ’### Subtask1: subtask1’.
Listing 12: Prompts for executing subtasks in action “Decompose”
1This is the problem context and the steps taken.
2{Problem and reasoning steps before decomposition}
3We have divided the main task into several parts. Below is the reasoning progress for the initial parts.
4{Executing results of previous subtasks}
5Now, please execute the specific instruction for Subtask{subtask_id} to advance the solution.
6DONOT conduct a more detailed decomposition for the subtask.
Listing 13: Prompts for summarize subtasks in action “Decompose”
1Below are the detailed execution steps for the problems subtasks.
2{Executing results of subtasks}
3Synthesize the steps above into a brief and coherent overview, ensuring that the essential logic and final outcomes of each subtask are preserved.
Debate: We rephrase the prompts for the action of Debate as:
Listing 14: Prompts for obtaining various plans in action “Debate”
1This is a problem and existing reasoning steps.
2{Problem and previous reasoning steps}
3Formulate three distinct strategies to address the problem from the current state.
5Use the format ’### Plan1: plan1’.
Listing 15: Prompts for analysing and comparing plans in action “Debate”
1This is the problem context and the steps taken.
2{Problem and previous reasoning steps}
3We have generated a set of potential strategies to address the problem at this stage. {Generated Plans}
4Critically evaluate these options and identify the most viable strategy to advance the reasoning.
5Only compare the plans here, and DONOT conduct further reasoning or calculation.
Listing 16: Prompts for executing the plan in action “Debate”
1Here is a problem and several reasoning steps.
2{Problem and previous reasoning steps}
3We have selected the optimal strategy to proceed: {Plan}
4Carry out the immediate next logical step following this strategy, and DONOT reason too many steps at once.
Refine: We rephrase the prompts for the action of Refine as:
Listing 17: Prompts for action “Refine”
1Here is a problem and several reasoning steps
2{Problem and previous reasoning steps}
3Please examine the current reasoning for potential improvements and revise it accordingly, and DONOT conduct further reasoning or calculation.
Table 23: Impact of prompt rephrasing.
LLM Method GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
RLoT (original) 92.87 44.64 79.04 72.18
RLoT (rephrased) 92.49 45.09 79.91 72.42
The results, presented in Table 23, demonstrate that RLoT exhibits high robustness to prompt variations. The model with rephrased prompts achieves an average accuracy very similar to the original one. This consistency indicates that the effectiveness of RLoT stems from the adaptive logical structures constructed by the navigator, rather than overfitting to specific phrasing or reliance on extensive prompt engineering.